# FairGU: Fairness-aware Graph Unlearning in Social Network

### v1

@@ 1. Introduction
As web applications continue to evolve into complex, interconnected ecosystems, graph-structured data has become the foundation of many large-scale online services (Zhang et al., 2024a), including social media (Vombatkere et al., 2024) and E-Commerce (He et al., 2024). In these settings, removing some privacy-related information from trained graph models is urgent (Li et al., 2025). Graph unlearning has emerged as a critical capability for protecting user autonomy and supporting ethical data governance,  (Fan et al., 2025). By enabling a model to remove a user’s data and its downstream influence, graph unlearning directly supports core values of an inclusive and sustainable web (Chen et al., 2025). This allows individuals to revoke consent while ensuring that platforms remain compliant with emerging privacy regulations and societal expectations, which is the right to be forgotten (Regulation, 2018). Such mechanisms are essential for building long-term trust between web platforms and their communities, a key requirement for achieving the broader vision of a web that serves the good of all.
@@
Existing graph unlearning methods are primarily designed to efficiently remove explicit personal information, including sensitive attributes, from the graph. However, a critical question remains: do these methods preserve downstream fairness associated with those attributes? In real-world web applications, algorithmic decisions must strictly adhere to principles of non-discrimination and maintain independence from protected attributes (Luo et al., 2024b). Algorithmic fairness thus provides a principled approach for evaluating whether a system treats users equitably, measured by independence criteria such as statistical parity (measured by $\Delta_{\text{SP}}$) and equality of opportunity (measured by $\Delta_{\text{EO}}$(Dong et al., 2023).
@@ 2.2. Fairness-aware GNNs
Parallel to developments in graph unlearning, significant research efforts have been dedicated to enhancing algorithmic fairness in GNNs, aiming to mitigate discriminatory predictions against sensitive subgroups defined by attributes such as race, gender, or region (Zhang et al., 2024b; Luo et al., 2024a). These methods primarily operate through pre-processing the data or in-processing during model training (Dong et al., 2023). Pre-processing methods like FairVGNN (Wang et al., 2022) and Graphair (Ling et al., 2023) aim to rectify bias at the data level by generating fair graph views or augmentations that weaken the correlation between sensitive attributes and other features. In scenarios with missing features, FairAC (Guo et al., 2023) jointly addresses attribute completion and unfairness, employing an attention mechanism to produce equitable node embeddings. Conversely, in-processing methods integrate fairness constraints directly into the learning objective. A prominent approach employs adversarial debiasing, as seen in FairGNN (Dai and Wang, 2023), which uses a sensitive attribute estimator and an adversarial loss to learn invariant node representations. Other methods modify the GNN architecture itself; for instance, FairSIN (Yang et al., 2024) introduces a neutralization-based paradigm by incorporating fairness-facilitating features during message-passing, while FMP (Jiang et al., 2024) explicitly restricts the use of sensitive attributes in forward propagation.
A common issue in many fairness-aware GNNs is their need to generate or estimate missing sensitive attributes (Luo et al., 2025). This approach can help reduce bias, but it also raises privacy concerns because it reconstructs information that users may not have wanted to reveal (Xiang et al., 2025). More critically from the perspective of this work, these models are designed for standard training scenarios and are not inherently equipped to handle the specific challenges of data removal requests (Zhang et al., 2025). When retrofitted for unlearning (e.g., using frameworks like IDEA (Dong et al., 2024) as a baseline adaptation), they often struggle to maintain fairness guarantees post-unlearning. This is because their debiasing mechanisms are optimized for the original data distribution and can be disrupted by the removal of nodes, potentially reintroducing or amplifying biases.
@@ 5.2. Baselines
Method Pokec-n Pokec-z Income
5.3. Experimental Setup
5.4. Comparison Results
@@ 5.5. Ablation Study
Table 4. Ablation study results of FairGU variants across datasets
@@ 6. Conclusion
In this paper, we propose FairGU, a new fairness-aware graph unlearning framework. It tackles the key challenge of applying fairness constraints during the data removal process in graph neural networks, especially when sensitive attributes are missing. FairGU ensures that after unlearning, the model removes the influence of the specified data while still maintaining fairness with respect to sensitive attributes. It achieves this through three components: a pre-trained sensitive-attribute estimator, a fairness-aware GNN with adversarial debiasing, and a precise unlearning mechanism built on Fisher Information Matrix analysis Extensive experiments on real-world datasets show that FairGU clearly outperforms SOTA methods. It achieves large gains in fairness metrics while still maintaining strong accuracy and providing robust privacy protection against membership inference attacks. Despite these strengths, FairGU’s performance is sensitive to hyperparameters such as the selectivity factor and dampening strength, which may require careful tuning in practical applications. Future work will focus on enhancing scalability and robustness, exploring adaptive hyperparameter optimization, and extending the framework to dynamic graphs and multi-sensitive attribute settings.

### v2

@@ 1. Introduction
As web applications continue to evolve into complex, interconnected ecosystems, graph-structured data has become the foundation of many large-scale online services (Zhang et al., 2024a), including social media (Vombatkere et al., 2024; Du et al., 2025) and e-commerce (He et al., 2024). In these settings, removing some privacy-related information from trained graph models is urgent (Li et al., 2025). Graph unlearning has emerged as a critical capability for protecting user autonomy and supporting ethical data governance (Fan et al., 2025). By enabling a model to remove a user’s data and its downstream influence, graph unlearning directly supports core values of an inclusive and sustainable web (Chen et al., 2025). This allows individuals to revoke consent while ensuring that platforms remain compliant with emerging privacy regulations and societal expectations, which is the right to be forgotten (Regulation, 2018; Song et al., 2025). Such mechanisms are essential for building long-term trust between web platforms and their communities, a key requirement for achieving the broader vision of a web that serves the good of all.
@@
Existing graph unlearning methods are primarily designed to efficiently remove explicit personal information, including sensitive attributes, from the graph. However, a critical question remains: do these methods preserve downstream fairness associated with those attributes? In real-world web applications, algorithmic decisions must strictly adhere to principles of non-discrimination and maintain independence from protected attributes (Luo et al., 2024b; Xu et al., 2022). Algorithmic fairness thus provides a principled approach for evaluating whether a system treats users equitably, measured by independence criteria such as statistical parity (measured by $\Delta_{\text{SP}}$) and equality of opportunity (measured by $\Delta_{\text{EO}}$(Dong et al., 2023; Oldfield et al., 2025).
@@ 2.2. Fairness-aware GNNs
Parallel to developments in graph unlearning, significant research efforts have been dedicated to enhancing algorithmic fairness in GNNs, aiming to mitigate discriminatory predictions against sensitive subgroups defined by attributes such as race, gender, or region (Zhang et al., 2024b; Luo et al., 2024a; Xu et al., 2025; Luo et al., 2025b). These methods primarily operate through pre-processing the data or in-processing during model training (Dong et al., 2023). Pre-processing methods like FairVGNN (Wang et al., 2022) and Graphair (Ling et al., 2023) aim to rectify bias at the data level by generating fair graph views or augmentations that weaken the correlation between sensitive attributes and other features. In scenarios with missing features, FairAC (Guo et al., 2023) jointly addresses attribute completion and unfairness, employing an attention mechanism to produce equitable node embeddings. Conversely, in-processing methods integrate fairness constraints directly into the learning objective. A prominent approach employs adversarial debiasing, as seen in FairGNN (Dai and Wang, 2023), which uses a sensitive attribute estimator and an adversarial loss to learn invariant node representations. Other methods modify the GNN architecture itself; for instance, FairSIN (Yang et al., 2024) introduces a neutralization-based paradigm by incorporating fairness-facilitating features during message-passing, while FMP (Jiang et al., 2024) explicitly restricts the use of sensitive attributes in forward propagation.
A common issue in many fairness-aware GNNs is their need to generate or estimate missing sensitive attributes (Luo et al., 2025a). This approach can help reduce bias, but it also raises privacy concerns because it reconstructs information that users may not have wanted to reveal (Xiang et al., 2025). More critically from the perspective of this work, these models are designed for standard training scenarios and are not inherently equipped to handle the specific challenges of data removal requests (Zhang et al., 2025). When retrofitted for unlearning (e.g., using frameworks like IDEA (Dong et al., 2024) as a baseline adaptation), they often struggle to maintain fairness guarantees post-unlearning. This is because their debiasing mechanisms are optimized for the original data distribution and can be disrupted by the removal of nodes, potentially reintroducing or amplifying biases.
@@ 5.5. Ablation Study
Table 4. Ablation study results of FairGU variants across datasets.
@@ 6. Conclusion
In this paper, we propose FairGU, a new fairness-aware graph unlearning framework. It tackles the key challenge of applying fairness constraints during the data removal process in graph neural networks, especially when sensitive attributes are missing. FairGU ensures that after unlearning, the model removes the influence of the specified data while still maintaining fairness with respect to sensitive attributes. It achieves this through three components: a pre-trained sensitive-attribute estimator, a fairness-aware GNN with adversarial debiasing, and a precise unlearning mechanism built on Fisher Information Matrix analysis Extensive experiments on real-world datasets show that FairGU clearly outperforms SOTA methods. It achieves large gains in fairness metrics while still maintaining strong accuracy and providing robust privacy protection against membership inference attacks. Despite these strengths, FairGU’s performance is sensitive to hyperparameters such as the selectivity factor and dampening strength, which may require careful tuning in practical applications.